State if the following statement is true or false. Give reasons for your answers in the form of a short proof or a counterexample.



(√2, 1, ½) E Q X Z X R.

1
Expert's answer
2022-06-08T18:03:10-0400

The statement is false because "\sqrt{2}\notin\mathbb{Q}".


Proof that "\sqrt{2}\notin\mathbb{Q}":

Suppose "\sqrt{2}" is rational. That mean it can be writen as the ratio of two inegers "p" and "q"

"\sqrt{2}=\frac{p}{q}" (1)

where we may assume that "p" and "q" have no common factors. (If there are any common factors we cancel tham in the numerator and denominator.) Squaring in (1) on both sides gives

"2=\frac{p^2}{q^2}" (2)

which impiels

"p^2=2q^2" (3)

Thus "p^2" is even. The only way this can be true is that "p" itself is even. But then "p^2" is actually divisible by "4". Hence "q^2" and therefore "q" must me even. So "p" and "q" are both even which is a contradiction to our assumption that they have no common factor. The square root of 2 cannot be rational.